Please explain how I may obtain the needed answer.
I have
(x+3)^2=-25
I thought
(X+3)*(x+3)=X^2+6x+9= -25 and moved into quadratic form by adding 25 to both sides. like so,
x^2+6x+34=0 then solve for X.
At this point I usually give the brain a break and solve using the TI89.
Problem is that it keeps telling me false instead of the needed, {-3+5i,-3-5i}
Can anyone tell me where I went wrong?
Thanks

Answer by stanbon(75887)

The sum of squares factors as follows:
a^2 + b^2 = (a+bi)(a-bi)
=================================
(x+3)^2=-25
---
(x+3)^2 + 25 = 0
Factor:
[x+3+5i][x+3-5i] = 0
x = -3-5i or x = -3+5i
========================
Not sure what your TI89 is doing.
========================
Cheers,
Stan H.
============